# Advaith Sethuraman, 2505 Hayward St., Ann Arbor, MI 48109

### v1

@@ Abstract
Open-source benchmark datasets have been a critical component for advancing machine learning for robot perception in terrestrial applications. Benchmark datasets enable the widespread development of state-of-the-art machine learning methods, which require large datasets for training, validation, and thorough comparison to competing approaches. Underwater environments impose several operational challenges that hinder efforts to collect large benchmark datasets for marine robot perception. Furthermore, a low abundance of targets of interest relative to the size of the search space leads to increased time and cost required to collect useful datasets for a specific task. As a result, there is limited availability of labeled benchmark datasets for underwater applications. We present the AI4Shipwrecks dataset, which consists of 24 distinct shipwreck sites totaling 286 high-resolution labeled side scan sonar images to advance the state-of-the-art in autonomous sonar image understanding. We leverage the unique abundance of targets in Thunder Bay National Marine Sanctuary in Lake Huron, MI, to collect and compile a sonar imagery benchmark dataset through surveys with an autonomous underwater vehicle (AUV). We consulted with expert marine archaeologists for the labeling of robotically gathered data. We then leverage this dataset to perform benchmark experiments for comparison of state-of-the-art supervised segmentation methods, and we present insights on opportunities and open challenges for the field. The dataset and benchmarking tools will be released as an open-source benchmark dataset to spur innovation in machine learning for Great Lakes and ocean exploration. The dataset and accompanying software are available at https://umfieldrobotics.github.io/ai4shipwrecks/.
@@ 1 Introduction
It is estimated that over $3$ million undiscovered shipwrecks lie on the ocean floor Gonzalez et al. (2009). Locating these submerged archaeological sites enables research into important maritime assets of historical significance. However, searching over large areas and vast depths of the sea requires expensive and time-consuming surveys, which inhibits new discovery of shipwreck sites. Marine robotic platforms, including autonomous underwater vehicles (AUVs), have demonstrated potential to carry out efficient, cost-effective large-area surveys of marine environments returning hundreds of gigabytes worth of data. Still, the interpretation of sonar imagery to identify sites of interest requires manual expert review. This can take many months to complete, often requiring multiple surveys to verify potential new discoveries.
@@
Figure 1: Our AI4Shipwrecks dataset aims to accelerate the development of shipwreck detection algorithms for sonar data collected onboard autonomous systems.
Figure 2: Data acquisition, processing, and network inference pipeline using the Iver3 autonomous underwater vehicle.
@@
Automated processing of sonar data collected over large-area surveys has the potential to accelerate the discovery of new sites of interest. On land, machine learning has led to great advances in computer vision and robot perception tasks, including object detection, semantic segmentation, and scene understanding. State-of-the-art machine learning methods rely on labeled datasets for supervised training of neural networks to learn pixel-wise segmentation predictions. However, for underwater domains, there is limited availability of publicly available, labeled datasets for sonar data due to challenges, time, and expense associated with data collection Singh and Valdenegro-Toro (2021); Ochal et al (2020).
@@
In this article we present a new dataset, AI4Shipwrecks, and we present benchmarking results for segmentation of shipwrecks from side scan sonar (SSS) imagery collected on an AUV (Fig. 1). This dataset was collected over the span of 5 weeks in Thunder Bay National Marine Sanctuary (TBNMS) in Lake Huron, Michigan, USA. Spanning 4,300-square-miles, TBNMS contains almost 100 known shipwreck sites and over 100 undiscovered sites. We leverage the unique abundance of known shipwreck sites to curate a rich dataset for investigating the application of machine learning for this task. In this paper, we present details of dataset collection and preparation for easy indexing of our dataset by machine learning pipelines. To encourage further advances in segmentation for side scan sonar, we report extensive results on our dataset from modern, state-of-the-art segmentation models. Lastly, we include discussion on lessons learned from our field expeditions and experiments to provide insight on future challenges and opportunities for machine learning for processing sonar data. The resulting dataset and benchmarking tools will be made publicly available as a benchmark dataset for segmentation from sonar imagery to enable future research in machine learning for ocean exploration.
@@ 2.1 Object Segmentation in Sonar Imagery
Sonar is a popular perception sensor underwater due to its long operational range Lin et al. (2023). As a result, sonar image understanding algorithms are crucial for enabling autonomous underwater vehicles to operate in unstructured environments. Since many sonar images can be manipulated as single-channel images, sonar image understanding algorithms have a high overlap with object detection and segmentation algorithms from computer vision.
Recent work focuses on the application of data-driven deep neural networks for object detection and segmentation in sonar imagery. The majority of work in object detection for sonar imagery involves retraining or fine-tuning existing object detection algorithms on task-specific datasets of sonar imagery Einsidler et al (2018); Yang et al. (2022). The unique nature of SSS imagery has also motivated the development of specialized network architectures Burguera and Bonin-Font (2020); Yang et al. (2022).
@@ 2.2 Datasets
Large, labeled, and publicly available SSS datasets incur high costs of collection and require expert analysis for labeling. As a result, there are very few options for evaluating the performance of object detection algorithms.
There has been recent interest in deep learning datasets for forward looking sonars (FLS) Singh and Valdenegro-Toro (2021); Choi et al. (2021); Xie et al. (2022); Santos et al. (2022). Singh and Valdenegro-Toro (2021) present an FLS dataset of debris in marine environments, but FLS is shorter range and has distinct sensor geometry compared to side scan sonar. Furthermore, the dataset is object-centric and captured in a controlled lab environment.
Although there is an abundance of unlabeled side scan datasets, they are also not immediately useful for benchmarking computer vision algorithms Bernardi et al. (2020). To the best of our knowledge, our AI4Shipwrecks dataset is the first publicly available, large, labeled dataset of side scan sonar images in the marine autonomy community.
Figure 3: Map of survey sites in TBNMS. Callouts include example sonar data overlaid with ground truth labels. Color indicates sites that are included in testing (red) and training (yellow) splits, and locations of additional terrain surveys (green).
@@ 2.3 Synthetic Datasets for Sonar
Simulation can produce low-cost, diverse, and labeled datasets for sonar image understanding Liu et al. (2021); Cerqueira et al. (2020); Sethuraman et al. (2023); Lee et al. (2018). Many simulators use ray-tracing techniques to render arbitrary 3D meshes Sethuraman et al. (2023); Cerqueira et al. (2020), while others use style-transfer techniques to mimic the noise and sensor models of sonars Lee et al. (2018); Sung et al. (2019). Although simulated data can help mitigate training data scarcity, it is still desirable to evaluate the performance of vision algorithms on a real test dataset. AI4Shipwrecks addresses this gap by leveraging the natural diversity of shipwreck sites in TBNMS. We present a dataset representative of real shipwreck sites an AUV could encounter during autonomous surveys.
@@ 3 Technical Approach
Figure 2 provides an overview of the developed pipeline for acquiring sonar data for machine learning applications. First, data is collected through deploying an AUV for a large-area survey. The AUV carries out a pre-programmed survey mission to acquire sonar data. Once the data are returned, post-processing converts the raw sonar data format (.JSF) to standard image format (.PNG). The standard image format is further processed to be input into a deep neural network. The network outputs a prediction in the form of a binary per-pixel segmentation mask, which can be visualized as an overlay on the input sonar image.
@@ 3.1 Site Selection
Surveys were conducted in Thunder Bay National Marine Sanctuary (TBNMS) in Lake Huron, MI. Figure 3 shows shipwreck sites that were imaged in our surveys during four weeks over the course of two years. The abundance of known shipwreck targets in TBNMS was a crucial factor in our selection of this field site, as this enabled us to maximize the number of targets observed within a relatively constrained area and short timespan. These proposed survey regions were selected in coordination with scientists at TBNMS to cover a wide range of ship types, site relief, wreck characteristics, and water depth, prioritizing sites within a reasonable distance from the port of Alpena. This ensured that we could survey a maximum number of sites while still capturing variation across samples, providing a unique and valuable dataset for training machine learning methods. Furthermore, this large and diverse dataset allows us to validate and thoroughly evaluate the accuracy and generalizability of developed methods.
@@ 3.2 Data Collection Platform
Figure 5: a) SSS sensor model detailing sensor tilt angle $\theta_{t}$, first bottom return $p_{fbr}$, nadir gap, and ensonified point on object $p(P_{x},P_{y},P_{z})$. The SSS field of view is shown in blue. b). The AUV (in yellow) performs a survey with half swath width $\frac{\lambda}{2}$, leg width $\delta_{w}$, leg length $l$, and total survey width $w_{s}$. For each survey leg, the AUV dives to the depth from surface $d_{s}$ or height from bottom $h_{b}$ and resurfaces between legs to acquire a GPS lock.
The surveys were conducted using the Michigan Technological University (MTU) Iver3 AUV (pictured in Fig. 4). Each survey mission was conducted by pre-programming a route to survey target sites. The AUV was programmed to capture SSS data while driving at a constant velocity of $2.5$ knots, with a modular sensor payload that includes an EdgeTech $2205$ dual frequency $600$/$1,600$ kHz ultra-high resolution side scan sonar and 3D bathymetric system. The released imagery is produced from the low frequency sonar. An upgraded SBG Ellipse-A Inertial Measurement Unit (IMU) installed on the AUV’s bathymetric sensor greatly improves the quality of this data, providing $<3$ cm of resolution in both the horizontal and vertical planes. The AUV’s Teledyne RDI Explorer Doppler Velocity Log (DVL) and high-accuracy compass provide absolute positional accuracy of $0.6$ percent of the distance traveled underwater (CEP95959595). The AUV is also equipped with a forward-looking obstacle avoidance sonar. This combination of stability, highly accurate 3D autonomous positioning, obstacle avoidance, and ultra-high resolution sonar and camera makes the Iver3 platform optimal for this application.
@@ 3.3 Side Scan Sonar Sensor Model
A side scan sonar emits fan-shaped low and high-frequency acoustic pulses from two transducers aimed at the seabed floor. The signal travels through the water column and reflects off the terrain or other objects in the swath area before being received by the sensor. After a sonar chirp has been emitted, the sensor receives and bins the intensity of returns according to time-of-flight. On the horizontal axis, an SSS image is a histogram of return intensity at equally spaced intervals in time. Each histogram is accumulated in the vertical axis as the transducer moves to produce an image with two dimensions. The higher the echo signal intensity, the higher the pixel value in the resulting sonar image. Raw SSS imagery is single channel grayscale but can be viewed with various color palettes for improved visibility.
The sensor model is depicted in Fig. 5a and illustrates an acoustic beam that encounters an object at point $p(P_{x},P_{y},P_{z})$. SSS typically has two identical transducer arrays mounted symmetrically on either side of the AUV at a fixed tilt angle $\theta_{t}$. The first return from the sonar is called the first bottom return ($p_{fbr}$) and is used as an estimate of AUV altitude. There is typically a gap between the swath widths of both sides of the sensor resulting in a nadir gap with no coverage. This causes a black stripe down the center of a sonar image, as seen in Fig. 5b.
Although side scan sonar produces high resolution images of the environment, view-dependent shadowing effects, self-occlusion, material-dependent acoustic noise, and distortion make object detection a difficult task for both humans and automated algorithms. Figs. 6d-f illustrate common distortions and noise found in side scan sonar imagery.
Figure 6: Images of wrecks highlighting levels of the three major categories used to split the dataset. Fig. 6a-6c represent the shipwreck conditions, Fig. 6d-6f represent different sonar quality conditions, and Fig. 6g-6i represent the terrain types.

### v2

@@ Abstract
Open-source benchmark datasets have been a critical component for advancing machine learning for robot perception in terrestrial applications. Benchmark datasets enable the widespread development of state-of-the-art machine learning methods, which require large datasets for training, validation, and thorough comparison to competing approaches. Underwater environments impose several operational challenges that hinder efforts to collect large benchmark datasets for marine robot perception. Furthermore, a low abundance of targets of interest relative to the size of the search space leads to increased time and cost required to collect useful datasets for a specific task. As a result, there is limited availability of labeled benchmark datasets for underwater applications. We present the AI4Shipwrecks dataset, which consists of 28 distinct shipwrecks totaling 286 high-resolution labeled side scan sonar images to advance the state-of-the-art in autonomous sonar image understanding. We leverage the unique abundance of targets in Thunder Bay National Marine Sanctuary in Lake Huron, MI, to collect and compile a sonar imagery benchmark dataset through surveys with an autonomous underwater vehicle (AUV). We consulted with expert marine archaeologists for the labeling of robotically gathered data. We then leverage this dataset to perform benchmark experiments for comparison of state-of-the-art supervised segmentation methods, and we present insights on opportunities and open challenges for the field. The dataset and benchmarking tools will be released as an open-source benchmark dataset to spur innovation in machine learning for Great Lakes and ocean exploration. The dataset and accompanying software are available at https://umfieldrobotics.github.io/ai4shipwrecks/.
@@ 1 Introduction
It is estimated that over $3$ million undiscovered shipwrecks lie on the ocean floor (Gonzalez et al., 2009). Locating these submerged archaeological sites enables research into important maritime assets of historical significance. However, searching over large areas and vast depths of the sea requires expensive and time-consuming surveys, which inhibits new discovery of shipwreck sites. Marine robotic platforms, including autonomous underwater vehicles (AUVs), have demonstrated potential to carry out efficient, cost-effective large-area surveys of marine environments returning hundreds of gigabytes worth of data. Still, the interpretation of sonar imagery to identify sites of interest requires manual expert review. This can take many months to complete, often requiring multiple surveys to verify potential new discoveries.
@@
Figure 1: Our AI4Shipwrecks dataset aims to accelerate the development of shipwreck segmentation algorithms for sonar data collected onboard autonomous systems.
Table 1: Table comparing existing publicly available sonar datasets for object detection and segmentation (FLS denotes forward looking sonar, and SSS denotes side scan sonar). Note that our dataset is the only one to offer pixel-level segmentation labels for shipwrecks in a real-world environment.
Dataset Environment
Sonar
Type
# Total
Images
# Shipwreck
Images
Segmentation
Labels
Marine Debris FLS (2021) Indoor Tank FLS 1868 0
SeabedObjects-KLSG (2020) Real World SSS 1190 385
AURORA (2022) Real World SSS 2 0
SeabedObjects-KLSG-II (2022) Real World SSS 1296 487
Marine-PULSE (2023) Real World SSS 719 0
Burguera and Bonin-Font (2020) Real World SSS 10 0
AI4Shipwrecks (ours) (2024) Real World SSS 286 161
@@
Automated processing of sonar data collected over large-area surveys has the potential to accelerate the discovery of new sites of interest. On land, machine learning has led to great advances in computer vision and robot perception tasks, including object detection, semantic segmentation, and scene understanding. State-of-the-art machine learning methods rely on labeled datasets for supervised training of neural networks to learn pixel-wise segmentation predictions. However, for underwater domains, there is limited availability of public, labeled datasets for sonar data due to challenges, time, and expense associated with data collection (Singh and Valdenegro-Toro, 2021; Ochal et al., 2020).
@@
In this paper, we present a new dataset, AI4Shipwrecks, and we present benchmarking results for segmentation of shipwrecks from side scan sonar (SSS) imagery collected on an AUV (Fig. 1). The AI4Shipwrecks dataset contains 286 high-resolution side scan sonar images in .PNG format, with accompanying pixel-wise segmentation labels of shipwreck sites. This dataset was collected over the span of five weeks in Thunder Bay National Marine Sanctuary (TBNMS) in Lake Huron, Michigan, USA. Spanning 4300-square-miles, TBNMS contains almost 100 known shipwreck sites and over 100 undiscovered sites. We leverage the unique abundance of known shipwrecks to curate a rich dataset of 28 distinct shipwrecks for investigating the application of machine learning for this task. In this paper, we present details of dataset collection and preparation for easy indexing of our dataset by machine learning pipelines. To encourage further advances in segmentation for SSS, we report extensive results on our dataset from modern, state-of-the-art segmentation models. Lastly, we include discussion on lessons learned from our field expeditions and experiments to provide insight on future challenges and opportunities for machine learning for processing sonar data. The resulting dataset and benchmarking tools will be made publicly available as a benchmark dataset for segmentation from sonar imagery to enable future research in machine learning for ocean exploration.
Figure 2: Data acquisition, processing, and network inference pipeline using the Iver3 autonomous underwater vehicle. The yellow lines in Mission Planning denote the AUV’s trajectory, with blue segments occuring at survey depth.
@@ 2 Background
Sonar is a popular perception sensor underwater due to its long operational range (Lin et al., 2023). As a result, sonar image understanding algorithms have the potential to enable autonomous underwater vehicles to operate in unstructured environments.
2.1 SSS Imagery vs. RGB Imagery
Converting the raw acoustic data to sonar imagery makes it more easily viewable by humans. However, there are key phenomena found in SSS imagery that are not present in RGB imagery. SSS imagery primarily measures the intensity of sound returned at a given range. Although it may look similar to an optical birds-eye-view of the underwater terrain, the SSS sensor model is dissimilar to RGB camera models. First, the sonar’s beam pattern unevenly distributes energy towards the seafloor, leading to varying image intensities along the horizontal axis (Burguera and Bonin-Font, 2020). Next, SSS is a time-of-flight sensor that exhibits acoustic shadows due to occlusions (Burguera and Bonin-Font, 2020). Finally, the resolution of SSS imagery depends on several parameters such as the altitude of the vehicle, range, and grazing angle, all of which can change within a single image (Grządziel, 2023).
@@ 2.2 Object Segmentation in Sonar Imagery
Since many sonar images can be manipulated as single-channel images after post-processing, sonar image understanding algorithms have many similarities to techniques intended for RGB images from computer vision.
Recent work focuses on the application of data-driven deep neural networks for object detection and segmentation in sonar imagery. The majority of work in object detection for sonar imagery involves retraining or fine-tuning existing object detection algorithms on task-specific datasets of sonar imagery (Einsidler et al., 2018; Yang, Wang, et al., 2022). The unique waterfall data format of SSS data has also motivated the development of specialized network architectures that enable real-time inference on streamed SSS data (Burguera and Bonin-Font, 2020).
@@ 2.3 Datasets
Large, labeled, and publicly available SSS datasets incur high costs of collection and require expert analysis for labeling. As a result, there are fewer options for evaluating the performance of object detection and segmentation algorithms meant for sonars. A summary of existing public sonar datasets for machine learning applications is shown in Table 1.
There has been recent interest in deep learning datasets for forward looking sonars (FLS) (Singh and Valdenegro-Toro, 2021; Choi et al., 2021; Xie et al., 2022; Santos et al., 2022). Singh and Valdenegro-Toro present an FLS dataset of debris in marine environments, but FLS is shorter range and has distinct sensor geometry compared to SSS. Furthermore, the dataset is object-centric and captured in a controlled lab environment.
There are two recent SSS datasets that contain submerged objects, namely SeabedObjects-KLSG (Huo et al., 2020) and Marine-PULSE (Du et al., 2023). However, neither of these datasets contain pixel-wise semantic labels. Bernardi et al. present SSS and AUV data of terrain in the Greater Haig Fras Marine Conservation Zone, however the dataset does not contain any shipwrecks and also does not have semantic labels. Similarly, the SSS dataset released by Burguera and Bonin-Font (2020) focuses exclusively on seafloor terrain. While this dataset does include pixel-wise labels, the labels are only for terrain types: “sand,” “rock,” and “other.” There are no shipwrecks included in the surveys and the dataset only contains a total of ten waterfall images.
To the best of our knowledge, our AI4Shipwrecks dataset is the first publicly available dataset for shipwreck segmentation in SSS images in the marine autonomy community.
Figure 3: Map of survey sites in TBNMS, Lake Huron, MI. Callouts include example sonar data overlaid with ground truth labels. Color indicates sites that are included in testing (red) and training (yellow) splits, and locations of additional terrain surveys (green). Best viewed in color and zoomed in.
@@ 2.4 Synthetic Datasets for Sonar
Simulation can produce low-cost, diverse, and labeled datasets for sonar image understanding (Liu et al., 2021; Cerqueira et al., 2020; Sethuraman et al., 2023; Lee et al., 2018). Many simulators use ray-tracing techniques to render arbitrary 3D meshes (Sethuraman et al., 2023; Cerqueira et al., 2020), while others use style-transfer techniques to mimic the noise and sensor models of sonars (Lee et al., 2018; Sung et al., 2019). Methods that train solely on simulated data encounter a sim-to-real gap between simulated sonar imagery and real imagery, which leads to reduced performance when testing on real data (Sethuraman et al., 2022). Prior work has studied the sim-to-real gap and techniques for reducing it (Sethuraman et al., 2023). Although simulated data can help mitigate training data scarcity, it is still desirable to evaluate the performance of vision algorithms on a real test dataset for deployment in the real world. AI4Shipwrecks addresses this gap by leveraging the natural diversity of shipwreck sites in TBNMS. We present a dataset representative of real shipwreck sites an AUV could encounter during autonomous surveys. We also release pixel-wise segmentation labels of shipwreck sites to enable thorough evaluation of machine learning methods for the task of shipwreck segmentation in real environments.
@@ 3 Technical Approach
Figure 2 provides an overview of the developed pipeline for acquiring and testing sonar data for machine learning applications. First, data is collected through deploying an AUV for a large-area survey. The AUV carries out a pre-programmed survey mission to acquire sonar data. Once the data are returned, post-processing converts the raw sonar data format (.JSF) to standard image format (.PNG). The standard image format is further processed to be input into a deep neural network. The network outputs a prediction in the form of a binary per-pixel segmentation mask, which can be visualized as an overlay on the input sonar image.
@@ 3.1 Site Selection
Surveys were conducted in TBNMS in Lake Huron, MI. Figure 3 shows shipwreck sites that were imaged in our surveys during five weeks over the course of two years. The abundance of known shipwreck targets in TBNMS was a crucial factor in our selection of this field site, as this enabled us to maximize the number of targets observed within a relatively constrained area and short timespan. We timed underwater surveys during late-May through early-June to mitigate the effects of thermocline. We did not conduct surveys in inclement weather. These proposed survey regions were selected in coordination with scientists at TBNMS to cover a wide range of ship types, site relief, wreck characteristics, and water depth, prioritizing sites within a reasonable distance from the Port of Alpena. This ensured that we could survey a maximum number of sites while still capturing variation across samples, providing a unique and valuable dataset for training machine learning methods. Furthermore, this large and diverse dataset allows us to validate and thoroughly evaluate the accuracy and generalizability of developed methods.
@@ 3.2 Data Collection Platform
The surveys were conducted using the Michigan Technological University (MTU) Iver3 AUV equipped with an EdgeTech $2205$ dual frequency $540/1610$ kHz SSS and 3D bathymetric system (pictured in Fig. 4). Note that the EdgeTech 2205 uses CHIRP and not continuous wave pulses (CW), and as a result there is a CHIRP start frequency ($C_{s}$) and CHIRP end frequency ($C_{e}$). Relevant sensor parameters are reported in Table 2. Each survey mission was conducted by pre-programming a route to survey target sites. The AUV was programmed to capture SSS data while driving at a constant velocity of $2.5$ knots. The released imagery is produced from the low frequency sonar, as a lower frequency is able to cover a wider swath of area compared to high frequency sonar. The AUV’s Teledyne RDI Explorer Doppler Velocity Log (DVL) provides a long-term accuracy of ±$\pm 0.3\%\pm 0.2$cm/s (Teledyne, 2024). The AUV’s DVL has a maximum range of 81 meters (Teledyne, 2024). The speed of sound in water is measured by an AML Oceanographic Sound Velocity Sensor, with a range of 1375-1625 m/s, precision of $\pm$ 0.006 m/s, accuracy of $\pm$ 0.025 m/s, and resolution of 0.001 m/s. The AUV is also equipped with a forward-looking obstacle avoidance sonar.
Table 2: Table with side scan sonar parameters. Note sonar range indicates the selected ranges during our surveys, not the min/max ranges for the sensor. The beam width $\theta_{h}$ is two-way.
Sensor Parameter Value
Center Frequency ($f$) 520 kHz
CHIRP Start Freq. ($C_{s}$) 488.5 kHz
CHIRP End Freq. ($C_{e}$) 551.5 kHz
Sweep Length ($\tau$) 1 ms
Horizontal Beam Width ($\theta_{h}$) 0.26 degrees
Survey Speed $(V_{s})$ 2.5 knots
Sonar Ranges $(R)$ 30-150 m
DVL Maximum Range 81 m
DVL Long Term Accuracy ±$\pm 0.3\%\pm 0.2$ cm/s
Sound Velocity Sensor Range 1375-1625 m/s
Sound Velocity Sensor Precision $\pm$ 0.006 m/s
Sound Velocity Sensor Accuracy $\pm$ 0.025 m/s
Figure 5: a) SSS sensor model detailing sensor tilt angle ($\theta_{t}$), first bottom return ($p_{fbr}$), nadir gap, and ensonified point on object ($p(P_{x},P_{y},P_{z})$). The SSS field of view is shown in blue. b) The AUV (in yellow) performs a survey with half-swath width ($\frac{s}{2}$), leg width ($\delta_{w}$), leg length ($l$), and total survey width ($w_{s}$). For each survey leg, the AUV dives to the depth from surface ($d_{s}$) or height over ground ($h_{g}$) and resurfaces between legs to acquire a GPS update. The SSS is only collecting data once submerged at depth ($d_{s}$) for length ($l$).
@@ 3.3 Side Scan Sonar Sensor Model
The EdgeTech 2205 SSS used in our surveys emits low and high frequency acoustic CHIRPs from two transducers aimed at the seafloor. The signal travels through the water column and reflects off the terrain or other objects in the swath area before being received by the sensor. After a sonar CHIRP has been emitted, the sensor receives and bins the intensity of returns according to time-of-flight. On the horizontal axis (across-track), an SSS image is a histogram of return intensity at equally spaced intervals in time. Each histogram is accumulated in the vertical axis (along-track) as the transducer moves to produce an image with two dimensions. The higher the returned signal intensity, the higher the pixel value in the resulting sonar image. Raw SSS imagery is single channel grayscale but can be viewed with various color palettes for improved visibility.
The sensor model is depicted in Fig. 5a and illustrates an acoustic beam that encounters an object at point $p(P_{x},P_{y},P_{z})$. SSS typically has two identical transducer arrays mounted symmetrically on either side of the AUV at a fixed tilt angle, $\theta_{t}$. The first return from the sonar is called the first bottom return ($p_{fbr}$) and is used as an estimate of AUV altitude. There is typically a sonar deadzone called a nadir due to the transducer beam pattern and geometry of the mounted sensor on the AUV. This causes a black stripe down the center of sonar images, as seen in Fig. 6.
Although SSS can produce detailed images of the environment, view-dependent shadowing effects, self-occlusion, material-dependent acoustic noise, and distortion make object detection a difficult task for both humans and automated algorithms. Figures 6d-f illustrate common distortions and noise found in SSS imagery.
3.4 Side Scan Sonar Resolution
SSS resolution is dependent on a variety of sensor and environmental parameters (Grządziel, 2023). These include the horizontal beamwidth ($\theta_{h}$), sonar range ($R$) reported in Table 2, and the speed of sound ($c$). Along-track resolution ($R_{x}$) is defined as the resolution in the direction of vehicle travel (vertical image axis):
$R_{x}=R\times sin(\theta_{h})$ (1)
Similarly, across-track resolution ($R_{y}$) refers to the resolution along the range axis (horizontal image axis). This is calculated based on the bandwidth (BW = $C_{s}-C_{e}$) of the sonar CHIRP used.
$R_{y}=\frac{c}{2\times BW}$ (2)
Using these formulas, we can calculate the resolutions of sonar imagery captured. We report along-track resolutions at the maximum sonar range and across-track resolutions at the measured speed of sound ($c$) in Table 3.
Figure 6: Images of wrecks highlighting levels of the three major categories used to split the dataset. Figures 6a-6c represent the shipwreck conditions, Figs. 6d-6f represent different sonar quality conditions, and Figs. 6g-6i represent the terrain types.